BEYOND THE NUMBERS 3.10 LEARNING OUTCOMES 1 TO 3 Bayes Rule EXTEND Name Section Number To be graded, all assignments must be completed and submitted on the original book page Background The notation P(A|B) is read as the probability of A, given B, has occurred. So the I symbol is read as given. Formally, A and B are called events and P(AIB) is a conditional probability Bayes rule is a very useful way of relating conditional and unconditional probabilities. According to this rule, for any two events A and B, we have P(B) Lets use T+ to denote the event the screening test concludes that the condition (disease, pregnancy, etc.) is present. Likewise T- is notation for the event the screening test concludes that the condition is not present. CP denotes that a condition is actually present, while CA denotes that a condition is actually absent. You already know some of the following definitions, though the notation is probably new to you. Others are likely all new: 1. P(T+ICP) is the sensitivity of the test 2, PT-ICA) is the specificity of the test. P(CPIT+) is the positive predictive value of the test. 4. P(CAIT- is the negative predictive value of the test. 5. P(CP) is the prevalence of the condition in the population. 3. EXHIBIT 1 Ottawa Ankle Rules Sensitivity of the Ottawa Rules G. Lucchesi, R., Jackson RE, W. Peacock WF, C. Cerasani, and R. Swor Title: Authors: Sor: Ann Emerg Med. 1995 Jul;26(1):1-5 The Ottawa ankle rules, commonly used in medicine, are designed to exclude fractures of the ankle (or midfoot). The rules first require a careful examination of two particular places on the ankle. If there is tenderness in either of those two places, or an inability to bear weight on the foot, then the Ottawa rules conclude that there is a fracture. Else, the rules conclude that there is not a fracture. In this publication, the authors conducted one of the first sensitivity and specificity studies on the Ottawa ankle rules. The following table did not appear in the referenced article, but was reconstructed from summary values that were recorded therein. A total of 421 patients with ankle injuries were studied. The truth regarding their fractures was confirmed with x-rays after the Ottawa ankle rules were applied 249 BEYOND THE NUMBERS 3.10Questions TABLE 3.10 Ankle Test Results Let CP represent the event that the ankle was truly fractured, and let T+ represent the event that the test said it was. From the table, estimate the positive predictive value and the negative predictive value 1. Truth Regarding Fracture Predicted by the Ottawa Ankle Test Totals Not Fractured Fractured 56 365 421 Not Fractured 51 Fractured 277 Totals 328 93 2. Which do you think a patient would be more interested in, P(T+ICP) or P(CPIT+)? Why? Have a look at the Bayes Rule equation again. Write out the equation for the positive predictive value. Assuming P(T+) doesnt change with prevalence, what happens to the positive predictive value as the prevalence in the population increases? Do you find that reasonable or not? Explain. 3. Do you think the Ottawa ankle rules are useful in the medical field? Why or why not? How do your analyses affect your position? 4. 1+ does, of course, change with prevalence, and this assumption is not necessary but sim not a required assumption. ies the question Your instructor may want to e lain·n more detail why it is 250 BEYOND THE NUMBERS 3.10
Answer #1

1) PPV = P(CP | T+) = P(CP and T+)/P(T+) = 88/365=   0.24109589          
              
NPV=P(CA|T-)=51/56=   0.910714286  

2)a patient would be more interested in P(T+|CP) i.e sensitivity of test

sensitivity represent the true positives.

  
3) PPV = P(CP | T+) = P(CP and T+)/P(T+)              
              
if T+ does not change with prevalence then PPV value does not change as prevalnce increases              
              
this does not seem reasonable because this will result in increase in false negatives              
              
4)sensitivity of test=88/93=94.62%

SPECIFICITY=51/328=15.55%

No, this is not useful,because specificity  is very less
              
true negatives are only 15.55%